### v2

# arXiv:2507.19364v2[cs.AI] 27 Feb 2026

Abstract
This position paper examines the use of Large Language Models (LLMs) in social simulation, analyzing their potential and limitations from a computational social science perspective. We first review recent findings on LLMs’ ability to replicate key aspects of human cognition, including Theory of Mind reasoning and social inference, while identifying persistent limitations such as cognitive biases, lack of grounded understanding, and behavioral inconsistencies. We then survey emerging applications of LLMs in multi-agent simulation frameworks, examining system architectures, scalability, and validation strategies. Projects such as Generative Agents (Smallville) and AgentSociety are analyzed with respect to their empirical grounding and methodological design. Particular attention is given to the challenges of behavioral fidelity, calibration, and reproducibility in large-scale LLM-driven simulations. Finally, we distinguish between contexts where LLM-based agents provide operational value—such as interactive simulations and serious games—and contexts where their use raises epistemic concerns, particularly in explanatory or predictive modeling. We argue that hybrid approaches integrating LLMs into established agent-based modeling platforms such as GAMA and NetLogo may offer a promising compromise between expressive flexibility and analytical transparency. Building on this analysis, we outline a conceptual research direction termed Hybrid Constitutional Architectures, which proposes a stratified integration of classical agent-based models (ABMs), small language models (SLMs), and LLMs within established platforms such as GAMA and NetLogo.
1 Introduction
Large Language Models (LLMs) have emerged as a transformative class of deep learning architectures capable of generating human-like text across a wide range of domains. Built upon the transformer architecture, which employs parallelized attention mechanisms rather than sequential processing, these models are pre-trained on massive corpora comprising web data, encyclopedic content, and code repositories [74, 20, 11]. Through unsupervised or self-supervised learning, LLMs acquire syntactic, semantic, and pragmatic language competencies, enabling them to perform a broad array of downstream tasks such as question answering, summarization, translation, and code generation [73]. Importantly, the magnitude and diversity of these training corpora mean that LLMs have effectively absorbed an immense statistical imprint of human language use across cultures, domains, and contexts. This goes far beyond learning grammar or syntax: LLMs encode traces of how people reason, argue, empathize, hesitate, and make decisions — all embedded in the linguistic fabric of their training data. Subsequently, these models are fine-tuned via instruction tuning, a supervised process in which models learn to follow natural language commands across a wide range of tasks. This is typically achieved using curated instruction datasets [58, 79], sometimes augmented with human feedback to better align model outputs with user intent [53].
Given this capability to emulate human-like reasoning patterns, LLMs have naturally attracted significant attention within the Agent-Based Modeling (ABM) community. However, as their integration into this field accelerates, it is crucial to distinguish between two emerging uses of Large Language Models in computational social science. The first involves using LLMs as coding assistants to facilitate the technical implementation of simulation models (e.g., generating NetLogo or GAMA code). The second, which is the sole focus of this paper, is the use of LLMs as the cognitive architecture of the agents themselves—often termed ’Generative Agents.’ We examine how LLMs can replace or augment traditional rule-based cognition to simulate human-like decision-making, social interaction, and reasoning, rather than how they assist in the software engineering process.
Focusing specifically on this role as a cognitive engine, a critical area of current investigation concerns the capacity of LLMs to authentically simulate human behavior. One classical benchmark in this domain is the Turing Test, proposed by Alan Turing in 1950 as the imitation game, which evaluates whether a machine can produce conversational behavior indistinguishable from that of a human interlocutor [35]. Recent empirical findings suggest that state-of-the-art LLMs, notably LLaMa-3.1 and GPT-4.5, have demonstrated convincing performance on a three-party version of this test. In these evaluations, LLaMa-3.1 was perceived as human in 56% of cases, while GPT-4.5 achieved a 73% success rate [35]. This represents a notable advancement over previous models, which had only succeeded under less rigorous two-party test conditions. These results underscore the impressive ability of modern LLMs to adopt complex linguistic styles and behavioral nuances, particularly when guided by carefully crafted persona prompts — a form of cognitive prompt engineering that remains a highly precise and non-trivial technique requiring expert design and task sensitivity.
The apparent success of LLMs in the Turing Test has naturally prompted speculation about their suitability for simulating human agents within computational social systems [54, 4, 39]. However, the capacity of LLMs to produce believable human-like dialogue is not indicative of genuine understanding or consciousness. Rather, their performance reflects advanced statistical pattern recognition — the ability to mimic linguistic structures convincingly — without necessarily modeling the underlying cognitive or emotional states of a human agent [35]. The tendency of human evaluators to anthropomorphize fluent language exacerbates this illusion, as observers may attribute intentionality or mental states to outputs that are, in essence, probabilistically generated [35], something well-known in psychology as the intentional stance [19]. This distinction between functional imitation and authentic comprehension is particularly salient for social simulation. If LLMs merely replicate surface-level features of human behavior, their utility for generating novel insights into social processes — rather than reproducing known patterns — is fundamentally limited. Moreover, their use as proxies for human agents in scientific studies raises questions about external validity, representational fidelity, and cognitive realism. These concerns constitute a central theme of this paper and inform our critical assessment of LLMs as generative agents in artificial societies.
This position paper proceeds as follows. Section 2 examines the psychological plausibility of LLMs as individual agents, evaluating their performance in social reasoning tasks while highlighting intrinsic cognitive limitations—such as hallucinations, inconsistency, and statistical biases—that challenge their reliability. Section 3 surveys the state of the art in LLM-based social simulation, reviewing emerging platforms, cognitive architectures for multi-agent systems, and current methodological strategies for empirical validation. Section 4 critically discusses the methodological barriers to scaling these agents into complex simulations, with particular emphasis on the ”micro-macro validity gap” the loss of epistemic control due to model opacity, and the ethical risks of automation bias. Finally, Section 5 proposes a synthesis of future research directions, examining the trade-offs between data-driven behavioral fine-tuning and theory-driven cognitive alignment. It advocates for ’Hybrid Constitutional Architectures’— a stratified paradigm that combines Small Language Models (SLMs) with theoretical psychological grounding—and discusses their integration into classical modeling environments (e.g., GAMA, NetLogo) to ensure groundedness and causal validity, particularly in data-scarce and culturally diverse contexts.
To ensure a comprehensive coverage of this rapidly evolving field, the literature selection for this position paper combines a systematic search with targeted exploratory retrieval. First, a structured query was conducted on the Scopus database (cutoff date: February 12) using the following boolean string: TITLE-ABS-KEY ((”Large Language Model*” OR ”Generative Agent*”) AND (”Social Simulation” OR ”Agent-Based”)). This initial search yielded a corpus of recent articles (2023-2026), from which we selected seminal works based on citation count and relevance to social science applications. Second, given the interdisciplinary and fast-paced nature of the domain (where many breakthroughs appear as preprints), we complemented this systematic extraction with LLM-assisted retrieval (using Gemini 3 Pro). This model were used to identify specific niche applications and to uncover cross-disciplinary links (e.g., cognitive psychology references) that keyword-based searches might miss. Crucially, human verification remained the final filter: every reference suggested by AI tools was manually reviewed by the authors to ensure scientific rigor, correct attribution, and contextual relevance to the arguments presented.
2 LLMs in Simulating Human Behavior: Capabilities and Cognitive Limitations
A critical question in assessing the suitability of LLMs for social simulation lies in evaluating how accurately they capture human psychological, emotional, and behavioral traits. Trained on massive corpora encompassing encyclopedic texts, social media exchanges, literature, forums, and dialogic interactions, LLMs have absorbed an unprecedentedly broad cross-section of human communicative behavior. As a result, they encode statistical traces of how people speak, reason, interact, and respond across a wide range of contexts — from mundane routines to emotionally charged situations. These models are thus not merely text generators, but synthesizers of large-scale sociolinguistic regularities, capable of emulating human-like behavior patterns precisely because such patterns are deeply embedded in their training data. Crucially, this capacity does not stem from intentional modeling or experiential understanding, but from the ability to compress and recompose behavioral signals at scale. When prompted appropriately, LLMs can reconstruct belief systems, emotional tones, and social scripts that mirror those present in the corpus — not through explicit design, but as emergent outcomes of large-scale pattern recognition. This gives rise to a powerful illusion: the model appears to “understand,” when in fact it is statistically interpolating between trillions of human-generated expressions.
Thus, while LLMs can convincingly mimic language use and pass superficial behavioral tests, the extent to which their outputs align with actual human cognition, reasoning patterns, emotional expression, and social bias remains contested.For instance, recent studies, such as [50], raise important concerns about the depth and consistency of LLMs’ psychological realism by exploring their potential to simulate stable and diverse personality traits. This section provides a structured review of the evidence regarding these capabilities and limitations.
2.1 Theory of Mind and Cognitive Reasoning Tasks
Theory of Mind (ToM) refers to the human capacity to attribute beliefs, intentions, desires, and knowledge to others. It is considered a foundational element of social cognition and typically emerges in early childhood. One of the canonical tests to assess ToM in developmental psychology is the Sally-Anne test [7]. In this task, a child observes a scene in which Sally places a marble in her basket and leaves the room. Anne then moves the marble to a box. Upon Sally’s return, the child is asked where she will look for her marble. A correct response—indicating Sally will look in the basket—demonstrates an understanding that Sally holds a false belief, which differs from the child’s own knowledge.
Recent studies have assessed the capacity of LLMs to pass such false-belief tasks. [40] tested multiple LLMs, including GPT-3.5 and GPT-4, on 40 ToM problems designed to mirror psychological benchmarks. While early models performed near chance levels, GPT-4 achieved over 75% accuracy, comparable to that of a 6 or 7-year-old child. Furthermore, GPT-4 demonstrated competence in second and third order ToM reasoning (e.g., “I believe that she thinks that he knows…”), with occasional success even at fifth and sixth-order levels, surpassing typical adult human performance in structured tasks.
The robustness of these findings has been explored in follow-up work. [32] tested LLMs on modified ToM tasks involving less prototypical narratives—such as a cat hiding a key or a user renaming folders offline—and found a significant drop in model performance. Performance was highly sensitive to phrasing; LLMs responded more accurately to questions using cognitive verbs (e.g., “think”) than to those phrased with action-oriented verbs (e.g., “look”), suggesting shallow linguistic pattern recognition rather than deep inference.
In addition to false-belief tasks, LLMs have been evaluated on their ability to interpret irony, indirect speech acts, and social narratives. [67] found that GPT-4 matched or exceeded average human performance in recognizing indirect requests, detecting sarcasm, and attributing communicative intent in narrative contexts. However, these results vary significantly with prompt design and context framing, raising questions about generalizability.
2.2 Emotion Representation and Affective Simulation
Emotions play a crucial role in human decision-making and social interaction. Unlike rule-based agents or earlier AI systems, LLMs can generate emotionally appropriate language in many contexts [45]. For example, given a prompt describing a tragic event, GPT-4 can produce empathetic responses and plausible emotional narratives. Nonetheless, this capacity is fundamentally linguistic. LLMs do not experience emotions nor maintain internal emotional states. Their affective responses are generated through statistical associations learned from text corpora containing human emotional expression.
Studies have attempted to evaluate LLMs’ performance on emotional recognition tasks. [42] report that LLMs can often classify emotions expressed in text (e.g., anger, sadness) and match tones to situational prompts. Yet these capabilities depend on typical lexical cues and falter when faced with indirect or culturally nuanced expressions. Emotional modeling in LLMs remains limited to surface-level mimicry, lacking integration with motivational systems or physiological grounding, both of which are central to human affective behavior.
2.3 Anthropomorphic and Statistical Biases
2.3.1 Societal and Representational Biases
Because LLMs are trained on large-scale human-generated data, they inevitably inherit and can perpetuate societal biases embedded in those texts. These include explicit and implicit stereotypes related to race, gender, religion, and culture. Studies have shown that LLMs reproduce associations consistent with biased representations: for instance, associating certain professions with male pronouns or producing racially skewed descriptions of criminal behavior [60]. In addition, structural imbalances in data availability mean that voices from underrepresented linguistic or social groups are poorly modeled, resulting in inequitable performance across different populations [41]. This ”black-box” inheritance of bias persists despite instruction tuning, and differences between model versions can lead to inconsistent user experiences and hidden harms in downstream applications [24].
2.3.2 Cognitive Biases and the Reasoning Paradox
Beyond representational biases, LLMs also exhibit patterns consistent with cognitive biases observed in humans—defined as systematic deviations from rationality in judgment. [38] found that LLMs replicate anchoring effects, status quo bias, framing effects, and recency bias when posed with structured reasoning tasks. While debiasing efforts using prompt engineering have shown promise, they remain partial solutions [62]. Alarmingly, the tendency to apply ”chain-of-thought” prompting—a method designed to enhance logical reasoning—does not consistently prevent these biases. In some instances, reasoning chains actually amplify them, indicating that LLM ”reasoning” may reinforce pre-trained associations rather than correct them [38]. [62] argues that mitigating these issues requires upstream control over training data composition rather than just refined prompting.
2.3.3 Convergence towards the ”Average Persona”
Another bias that raises significant ethical issues is the convergence towards an “average persona.” LLMs fundamentally operate on statistical probabilities, predicting the ”next likely word” based on patterns in their training data rather than deep cognitive understanding [75, 57]. Consequently, while human decisions are shaped by a rich tapestry of lived experiences and future goals, an LLM merely calculates the most mathematically expected response. This results in a flattening of diversity that can inadvertently erase the characteristics of minority subgroups in simulated populations [81, 24]. Empirical evidence highlights this homogeneity in economic games, where LLMs often produce highly uniform responses compared to the wider range of choices observed in humans [81]. This trend extends to the political sphere, where models tend to produce narrower distributions of opinions that overrepresent the values of wealthy, young, and politically liberal individuals—the so-called ”WEIRD” demographic bias [59, 29].
2.3.4 Artifactual Biases from Safety Alignment
Beyond this statistical convergence, the deliberate ”safety alignment” of LLMs—achieved through Reinforcement Learning from Human Feedback (RLHF)—introduces a further critical risk: the generation of ”false negatives” in exploratory research. While training models to be ”helpful and harmless” is essential for commercial products, this normative alignment acts as a systematic filter that ”sanitizes” agent behaviors. In the context of social simulation, an over-aligned model may exhibit a ”sycophancy bias” [63], refusing to manifest conflictual, aggressive, or ”dark” social dynamics even when such behaviors are realistic and essential for understanding phenomena like polarization or radicalization. Consequently, researchers risk exploring a truncated solution space where catastrophic but plausible social outcomes are suppressed, leading to a dangerous underestimation of systemic risks. While context-rich prompts can mitigate this convergence, this variability remains contingent on the user’s intervention rather than an inherent capacity of the model for autonomous diversity [5].
2.3.5 Bounded Rationality vs. Artifactual Bias
Finally, a critical theoretical distinction must be made regarding the very definition of ’bias’ in generative social science. Following Herbert Simon’s theory of Bounded Rationality [65], human decision-making is inherently constrained by cognitive limitations and heuristics. In social simulation, these human deviations from pure logic are not defects to be removed, but constitutive features necessary for realism. Therefore, we must rigorously distinguish between ’Desirable Cognitive Biases’—such as loss aversion or emotional reactivity, which reflect authentic human irrationality—and ’Artifactual Model Biases’ stemming from the LLM’s training process. The latter include phenomena like ’Sycophancy’ (the tendency to agree with the user to be ’helpful’) [63] or ’Safety Sanitization’ (the refusal to simulate conflictual interactions). While Simon’s bounded rationality is a target to be reached, these artifactual biases are ’machine-like’ distortions that flatten social dynamics and must be actively mitigated through architectural constraints.
2.4 Intrinsic Instability: Hallucinations and Inconsistency
A pervasive challenge in LLM behavior is the issue of hallucinations, where the model generates outputs that are baseless, untruthful, or diverge from factual reality, even while maintaining a coherent linguistic structure [73]. These errors often stem from flawed training data or specific architectural biases [48, 31, 28]. Crucially, when LLMs hallucinate, they typically maintain an ’aura of authority’, generating responses that appear coherent and well-formed. This deceptive coherence creates a ’validity illusion’ that can mislead users—especially those lacking domain expertise—and introduces substantial risks in scientific applications where factual accuracy is paramount [64, 68].
Compounding this issue is the problem of stochastic inconsistency. Identical or nearly identical inputs can elicit divergent responses, even within the same session, which hinders reproducibility and challenges the interpretability of agent behavior [23]. This instability can leave lasting impressions of unreliability, limiting the trustworthiness of the simulation [25]. Furthermore, while LLMs demonstrate superior performance in unrealistic, ’omniscient’ settings where all agents possess full information [1], they struggle significantly in scenarios characterizing real-world conditions, such as information asymmetry or the management of private knowledge [86]. Paradoxically, despite this individual instability, Wu et al. (2025) emphasize that LLM agents, when evaluated across multiple instantiations, often display lower variance and reduced behavioral diversity compared to human participants, amplifying social stereotypes and tending towards a ’mode collapse’ [81]
2.5 Summary and Implications
The current landscape presents a paradox. LLMs exhibit impressive capabilities in tasks traditionally considered strong indicators of social cognition. They can perform multi-level mental state inference, recognize irony, and produce emotionally plausible responses. However, these abilities are brittle, highly sensitive to phrasing, and grounded in probabilistic mimicry rather than experiential or embodied cognition. Their outputs reflect statistical associations from training corpora, not inferential reasoning rooted in a coherent world model.
In light of these findings, LLMs offer valuable opportunities for simulating linguistic and behavioral patterns within constrained contexts. Yet, their application in scientific investigations of human psychology or complex social behavior should be approached with caution. Their limitations—opacity, hallucination, inconsistency, bias, and shallow understanding—necessitate robust validation, controlled environments, and clear demarcation between imitation and cognition. In the following section, we turn to how these capabilities are being harnessed in practical social simulation frameworks, and how model architecture and evaluation are being addressed in emerging systems.
3 State of the Art: LLM-Based Social Simulation
The use of LLMs in social simulations builds upon years of research that have seen the application of machine learning techniques in multi-agent systems and, by extension, in agent-based models [22]. It is therefore essential to understand whether LLMs represent a major breakthrough or, more modestly, a logical evolution of this partnership, and to determine if they share the same methodological challenges as previous approaches or introduce new ones. Building on the previous discussion regarding the representational capacities of LLMs in psychological and behavioral contexts, this section examines their application in the field of social simulation. We focus on how they are currently used to instantiate artificial societies, the architectural strategies adopted for agent based modeling, and the methodological challenges related to validation and scaling.
3.1 From Generative Agents to Population-Scale Simulation
A foundational illustration of LLM-based social simulation is the Generative Agents framework introduced by [54], widely known as the “Smallville” project. In this system, twenty-five agents inhabit a sandbox environment reminiscent of The Sims, where they autonomously engage in daily routines, form interpersonal relationships, and generate emergent collective behaviors. Each agent maintains a natural language memory stream, allowing it to reflect on past events, summarize experiences, and formulate plans. Remarkably, the simple assignment of a goal—such as organizing a party—was sufficient to initiate a chain of spontaneous, coordinated interactions among agents. Evaluation efforts focused primarily on assessing the believability and internal coherence of these emergent dynamics, rather than on their correspondence to empirical data.
More recent efforts have extended this line of research to address questions of scale, complexity, and empirical grounding. One prominent example is AgentSociety, a large-scale platform designed to simulate the dynamics of human societies with over 10,000 LLM-based agents [56]. Each agent is instantiated using GPT-4 and equipped with cognitive modules for memory, goal management, decision-making, and social relationship tracking. These agents interact within rich environments encompassing spatial mobility, economic activity, social networks, and exogenous events such as natural disasters or political crises. AgentSociety has been applied to simulate phenomena including political polarization, rumor diffusion, and the societal impacts of universal basic income. Its validation strategy emphasizes empirical fidelity by replicating outcomes observed in behavioral experiments and large-scale social surveys.
Beyond general-purpose social interaction, emerging frameworks are now targeting specific socio-technical systems where physical constraints are paramount. Notably, GATSim (Generative Agent Transport Simulation) is a specialized cognitive framework for urban mobility simulation [49]. Moving away from traditional utility-based discrete choice models, their generative agents demonstrate the ability to plan complex, realistic daily travel itineraries. This work highlights a crucial evolution: the application of LLMs is shifting from purely conversational environments (like Smallville) to spatially constrained domains, where agents must balance personal preferences with physical transportation networks .
Complementing these large-scale efforts, Simulate Anything offers a flexible platform for generating demographically diverse agent populations using LLMs, leveraging real-world user data for initialization and sampling [78]. The system supports scenario-specific tuning and hypothesis testing across domains such as politics, journalism, and economics. Its principal contribution lies in enabling heterogeneous role configurations and behavioral variability, though its validation methodologies are comparatively underdeveloped.
Another significant contribution in this domain is S3 (Social-network Simulation System), which specifically focuses on social network dynamics through LLM-empowered agents [26]. S3 demonstrates particular strengths in simulating three pivotal aspects of human social behavior: emotion, attitude, and interaction behaviors. It leverages advanced prompt engineering and prompt tuning techniques to ensure authentic human-like agent behavior within social networks.
Several more recent platforms have further advanced the field in terms of architectural modularity and methodological rigor. GenSim, for instance, is a general-purpose simulation engine that supports tens of thousands of LLM agents [72]. It introduces structured abstractions for social routines, long-term memory management, and adaptive behavior correction, all orchestrated by a dynamic scheduling system. Validation combines empirical replication, face validity assessments, and anomaly detection to ensure both plausibility and statistical consistency.
Another emerging contribution is AgentTorch, a modular and scalable framework that unifies large-scale agent-based modeling with differentiable programming and LLM-powered behavior synthesis [15]. Instead of relying on expensive real-time prompting for each agent, AgentTorch uses a small number of archetypal LLM agents to generate compact, reusable policy representations. These are then deployed across massive populations of lightweight agents in GPU-accelerated simulations. Whereas narrative-first systems emphasize fine-grained agent individuality, AgentTorch focuses on experimental control, statistical robustness, and scientific reproducibility, making it particularly suited for large-scale hypothesis testing and parameter exploration in computational social science.
SALLMA (Scalable Architecture for LLM Multi-Agent Applications) adopts a layered architecture, separating operational processes (intent formation, task execution, communication) from knowledge-level components (agent profiles, shared memory, workflows) [8]. This modular design facilitates agent specialization and integration with external tools via APIs. Its validation pipeline combines automated benchmarking with expert-in-the-loop review, aided by interactive dashboards and simulation playback for iterative refinement.
SocioVerse represents a shift toward population-scale calibration [85]. It initializes agents from millions of real-world user profiles to mirror authentic distributions of attitudes, preferences, and social structures. The platform supports fine-grained tuning for scenario exploration, such as testing the effects of policy changes across diverse subpopulations. Its validation strategy emphasizes both empirical benchmarking and participatory evaluation, particularly focusing on the accurate representation of minority groups and on interactive adjustment of simulation parameters.
Similarly, the LLM-AIDSim framework integrates LLMs (Llama3:8b) into ABM for influence diffusion, enabling language-level responses and deeper insights into user agent interactions [84].
Collectively, these platforms reflect the rapid evolution of LLM-based social simulation toward greater scale, demographic realism, and methodological transparency. They also signal a growing reliance on hybrid validation strategies that combine quantitative benchmarking, expert evaluation, and dynamic, interactive diagnostics. Together, they mark a shift toward simulation systems that not only generate plausible behavior but also aspire to meet the standards of interpretability and scientific robustness necessary for credible computational social science.
3.2 Architecture Patterns for Cognitive Agents
LLM-based multi-agent models generally instantiate each agent as a distinct LLM instance or session using prompt engineering. Architectures typically involve three interconnected modules: a memory system that encodes and retrieves past experiences; a reflection and summarization layer that condenses observations into mental models; and a planning component that uses these reflections to generate actions [54, 46]. These modules are often implemented as chained or nested prompts, with each step leveraging the LLM’s generative capacity to simulate cognitive processes such as recall, reasoning, and intention formation. For instance, the Humanized Agent-Based Models (h-ABM) framework, proposed by [6], outlines a modular architecture for integrating LLMs into ABMs to achieve more cognitively and emotionally realistic agents.
3.2.1 Memory Systems and Psychological Foundations
Many LLM-based agent architectures draw inspiration from cognitive psychology, particularly theories of episodic and semantic memory. For example, the memory systems in Generative Agents and AgentSociety are designed to mimic human-like recall, using vector databases or embedding-based retrieval to surface relevant past experiences. Some frameworks explicitly reference psychological models, such as the Atkinson-Shiffrin multi-store model or the concept of “theory of mind,” to inform the design of agent cognition and social reasoning [46, 21]. Reflection modules often prompt the LLM to summarize recent events, extract intentions, or infer the mental states of other agents, thus operationalizing constructs from social and developmental psychology.
Established cognitive architectures such as SOAR [43] and ACT-R [3] have provided theoretical foundations for agent design. SOAR, developed by John Laird, Allen Newell, and Paul Rosenbloom at Carnegie Mellon University [43], embodies the problem space hypothesis, suggesting that all goal-oriented behavior can be characterized as a search in a problem space defined by states, operators, and goals. This cognitive architecture divides knowledge into procedural memory (skills and know-how), semantic memory (facts about the world), and episodic memory (memories of experiences). Similarly, ACT-R (Adaptive Control of Thought-Rational) [3] posits that human cognition can be modeled through interconnected cognitive and perceptual modules that operate according to rational principles. These architectures inform the contemporary design of LLM agents by providing structured frameworks for organizing memory, decision-making, and learning.
3.2.2 Planning and Decision-Making
Planning modules in LLM-based multi-agent systems typically exploit the model’s capacity to generate contextually appropriate actions by conditioning outputs on agent-specific goals, environmental affordances, and social context. While a subset of frameworks integrates reinforcement learning or utility-based decision mechanisms to refine agent strategies, the majority rely on prompt engineering and scenario-specific instructions to elicit desired planning behavior. This approach emphasizes the structured design of inputs to steer agents’ reasoning, allowing for complex decision-making without the need for explicit reward modeling.
A hybrid approach is exemplified by the LGC-MARL framework proposed by [33], which combines an LLM-based planner with graph-structured multi-agent reinforcement learning. In this architecture, the LLM is tasked with decomposing high-level instructions into executable subtasks, evaluating their coherence and feasibility, and formulating collaborative strategies among agents before final action selection. In contrast, more conventional social simulation systems rely predominantly on prompt-engineering strategies to define agent roles, goals, and reasoning pathways, typically without integrating reinforcement signals or dynamic adaptation mechanisms [36]. These approaches underscore the current design trade-off between architectural complexity and ease of interpretability in the planning components of LLM-driven simulations.
3.2.3 Orchestration and Communication
In LLM-based multi-agent systems, a centralized orchestration layer is commonly adopted to manage simulation time-stepping and regulate agent interaction flows. Such architectures typically support inter-agent communication via various message-passing mechanisms, including direct messaging, simulated surveys, and public broadcasts.
Approaches to orchestration and communication vary considerably depending on the complexity and goals of the simulation. For example, platforms like Generative Agents and AgentSociety utilize orchestration layers to manage complex interactions among thousands of agents, often based on sequential flows or perception-reflection-action loops. Others, such as GenSim, integrate structured abstractions for social routines and a dynamic scheduling system. More advanced frameworks, like AgentTorch, move away from costly sequential dependencies by generating reusable policy representations from a small number of archetypal LLM agents, which are then deployed at scale on lightweight agents for GPU-accelerated simulations. SALLMA proposes a layered architecture, separating operational processes (intention formation, task execution) from knowledge-level components (agent profiles, shared memory, workflows), thereby facilitating agent specialization and integration with external tools via APIs. These platforms often support modular ”blocks,” which are reusable behavioral units that allow for systematic experimentation and increased extensibility. Integration with external APIs (e.g., web search, databases, calculators) is also crucial, as it enables agents to access additional information and tools, thereby augmenting their reasoning and decision-making capabilities [14].
3.2.4 Specialization and Behavioral Diversity
LLM-based multi-agent system (MAS) frameworks are progressively integrating sophisticated specialization protocols, allowing individual agents to adopt roles characterized by unique behavioral repertoires. This specialization is typically realized through several mechanisms: the application of tailored prompt templates that guide an agent’s responses (e.g., [80]), which can enforce specific communication styles or decision-making heuristics; the implementation of role-specific memory structures that store relevant knowledge and past interactions (e.g., [54]), enabling agents to leverage historical data pertinent to their designated function; and scenario-driven fine-tuning processes that adapt an agent’s core model to specific operational contexts (e.g., [53]), thereby optimizing their performance for particular tasks or environments. Furthermore, some cutting-edge systems are exploring methods to inject personality traits or demographic attributes into agents. This can be achieved either by conditioning prompts with specific personality descriptors (e.g., ”act as a skeptical researcher”) or by sampling attributes from real-world survey distributions (e.g., age, occupation, cultural background), thereby enhancing the realism and diversity of agent behaviors and fostering more nuanced social dynamics [77, 14]. These innovative techniques clearly enhance existing population synthesis algorithms (e.g., [12]) and enable researchers to explore complex emergent behaviors within agent populations, thereby gaining deeper insights into human-like interactions.
3.3 Current Validation Methodologies
3.3.1 Empirical Validation and Benchmarking
The validation of LLM-based social simulations remains an active and evolving area of research. Among the most rigorous efforts, frameworks like AgentSociety adopt multi-faceted validation strategies that compare simulated outputs with real-world data. Empirical benchmarking typically involves aligning simulation results with social indicators such as voter turnout, economic behavior, or survey-derived attitudinal distributions from established instruments like the General Social Survey or the World Values Survey [56]. In addition, some studies replicate classical experimental paradigms—such as public goods games or rumor propagation scenarios—within the simulation environment, and evaluate whether the emergent outcomes converge toward established empirical findings.
Longitudinal consistency is also a critical focus, particularly in simulations involving extended time horizons. Researchers analyze whether agent behavior remains stable and coherent across simulation runs, and whether emergent patterns are reproducible under minor changes to initial conditions or prompt phrasing [81]. These approaches aim to assess not only face plausibility but also the robustness and generalizability of the simulated social dynamics.
3.3.2 Human-in-the-Loop Evaluation
Despite advances in automation, human judgment continues to play a crucial role in validating LLM-driven simulations. Expert evaluation by domain specialists is frequently used to assess the believability, internal coherence, and sociological plausibility of agent behaviors. Structured rating protocols are often employed to ensure inter-rater reliability, although such methods remain sensitive to subjective bias. Crowdsourced evaluation has also been adopted in some frameworks, wherein large groups of non-expert raters assess the realism or diversity of agent interactions.
Calibration against real-world survey data constitutes another widely used strategy. Here, demographic or attitudinal profiles from empirical sources are used to condition LLM agents, and human evaluators subsequently assess the extent to which the simulated agents reflect the intended population characteristics [14]. Nevertheless, recent findings suggest that human raters may be influenced by non-epistemic factors—such as assertiveness, linguistic fluency, or the perceived authority of LLM outputs—leading to potential overestimation of their factual or analytical accuracy [13].
3.3.3 Specialized Validation Methods
In response to the methodological challenges of evaluating LLM-generated behavior, several specialized validation approaches have emerged. Face validation techniques rely on the visualization of agent behavior, often through animation or dashboards, to elicit expert feedback on whether the behavior appears theoretically plausible [54]. Exploratory model behavior analysis is another technique, in which researchers systematically vary input conditions to test whether emergent behaviors align with theoretical expectations. Statistical correspondence testing has been used to quantitatively compare simulation outcomes to known empirical distributions or outputs from analytical models. Sensitivity analysis further complements these methods by evaluating the robustness of simulation results to perturbations in initial conditions and parameter values.
Sensitivity analysis further complements these methods by evaluating the robustness of simulation results to perturbations in initial conditions and parameter values. One concrete instantiation of this multi-level validation approach is the S3 system [26], which conducts evaluations at both individual and population levels using real-world social network data. Their validation strategy demonstrates promising accuracy in replicating observed social phenomena, including the propagation of information, attitudes, and emotions across network structures.
3.3.4 Data Sources and Comparative Benchmarks
The development of standardized benchmarks has been instrumental in advancing the empirical evaluation of LLM agents in social simulations. For instance, the SocialIQA dataset [61] evaluates an LLM’s capacity for common sense social reasoning by presenting situations involving physical, social, and emotional causality and testing whether the model can infer likely consequences. Another example is CrowS-Pairs [51], which assesses social bias by contrasting sentence pairs that differ only by protected attributes (e.g., gender or race), allowing systematic measurement of discriminatory tendencies. Additionally, benchmarks like BiosBias [18] evaluate demographic fairness by examining whether the model’s predictions or representations vary unjustly depending on demographic factors such as gender or ethnicity. While these datasets do not address agent consistency across sessions directly, related work such as [83] examines the consistency of language models by testing response variability to semantically equivalent paraphrased inputs, providing useful insights into reproducibility concerns in agent behavior.
Comparative studies between human and LLM behavior further inform the validation landscape. The EPITOME battery [37] includes a series of experimental tasks designed to test Theory of Mind capabilities in both humans and LLMs. Results suggest that while LLMs can often match or even exceed human-level performance on certain tasks, they also exhibit systematic limitations, particularly in pragmatic reasoning and inference under uncertainty.
Together, these empirical, human-centered, and dataset-driven approaches constitute an emerging ecosystem of validation methodologies. Nevertheless, their integration into a cohesive and principled framework remains an open area of research—one that is essential for ensuring the scientific credibility and social relevance of LLM-based simulations.
3.4 Implications and Methodological Cautions
Integrating LLMs into social simulation offers compelling opportunities: simulations are faster to build, more flexible in scope, and capable of generating richly expressive agent behavior. However, they also introduce epistemic risks. LLMs, trained to imitate plausible language rather than model causal processes, may yield emergent behaviors that are linguistically coherent but sociologically superficial or misleading. The gap between internal statistical inference and actual human cognition (as discussed in the previous section) persists in these settings and is amplified at scale. As a result, LLM-based social simulations may be most suitable for exploratory modeling, hypothesis generation, or interactive training environments—rather than confirmatory empirical research. Their use demands rigorous validation frameworks, careful scenario scoping, and, where appropriate, hybrid integration with traditional rule-based or equation-based models. Future work must focus on developing principled approaches for behavioral diversity injection, cross-run reproducibility, and modular interpretability to ensure these systems can meaningfully contribute to computational social science.
4 Methodological Challenges and Epistemological Risks
The transition from the technical frameworks described in the previous section to rigorous scientific inquiry reveals a distinct set of hurdles. While the platforms reviewed above demonstrate that large-scale LLM simulations are technically feasible, feasibility does not equate to scientific validity. The very features that make LLMs attractive—their ability to generate fluent, varied, and human-like text at scale—also introduce profound methodological barriers when moving from exploratory modeling to causal explanation. This section critically examines these barriers, focusing on the gap between micro-level believability and macro-level veridicality, the loss of epistemic control due to model opacity, and the systemic risks of bias and cost. To guide this analysis, Table 1 provides a synthetic overview of these cognitive, methodological, and technical hurdles, alongside the architectural solutions proposed to overcome them.
Table 1: Taxonomy of Challenges in LLM-based Social Simulation and Proposed Hybrid Mitigations. This table synthesizes the methodological barriers identified in Section 4 and maps them to the architectural solutions proposed in Section 5.
Category Challenge & Mechanism Impact on Social Simulation Hybrid Mitigation (Sec. 5)
Cognitive & Behavioral The “Average Persona” Risk
Convergence toward a statistical mean; suppression of heterogeneity [81].
Replicant Effect: Loss of population diversity; inability to model polarization or minority behaviors. Persona Injection & SLMs: Use specialized Small Language Models (SLMs) fine-tuned on local data rather than generic foundation models.
Generative Exaggeration
Tendency to caricature specific roles when prompted away from the mean [52].
Toxic Caricature: Agents exhibit stereotypical or extreme behaviors rather than nuanced human decision-making. Constitutional Filtering: Constrain LLM outputs using a theoretical framework (e.g., TPB) to reject implausible extremes.
Methodological & Epistemic The “Fluency Fallacy”
High linguistic coherence masking factual hallucinations [66].
Surface Realism: Simulations appear valid to human evaluators but lack causal veridicality. Groundedness via ABM: Use the ABM environment as the “ground truth” (physics, resources) to validate agent actions.
Data Scarcity Wall
Lack of training data for non-WEIRD populations [60].
Cultural Erasure: Inability to simulate specific groups (e.g., Vietnamese farmers) via pure fine-tuning. Structural Alignment: Replace missing data with theoretical rules (Constitutional Architecture) to guide reasoning.
Technical & Operational Temporal Instability
Context window limits and stochastic drift over long horizons [81].
Longitudinal Incoherence: Agents “forget” their history or change personality during long simulations. External Memory Modules: Offload long-term memory and state management to the classical ABM structure.
Computational Bottleneck
High inference costs preventing large-scale runs.
No Sensitivity Analysis: Inability to perform rigorous statistical validation (1000+ runs). Stratified Architecture: Delegate routine heuristics to lightweight SLMs; reserve LLMs for complex novelty.
4.1 The Validity Gap: From Micro-Believability to Macro-Veridicality
While frameworks for large-scale simulation have emerged, a critical methodological barrier remains regarding empirical validation. As argued by [44], validation is currently the central challenge for generative social simulation, as the field lacks standardized protocols to move beyond ’surface realism’. Most current validation strategies rely on ’micro-level believability’—assessing whether individual agents respond realistically in isolated contexts [54]. However, establishing micro-validity does not guarantee ’macro-level veridicality’.
This gap between individual plausibility and collective accuracy is widened by three systemic failures that only manifest at scale. First, the convergence toward an ”average persona” [81] suppresses the behavioral heterogeneity necessary for modeling complex social phenomena. In many domains, this results in a mode collapse where diverse agents regress to a uniform mean. Second, temporal instability—often exacerbated by context window limitations—compromises longitudinal coherence. Agents tend to ’forget’ early instructions or exhibit behavioral drift over long simulation runs, making reproducibility—a cornerstone of scientific inquiry—notoriously difficult [81]. Third, many simulations rely on omniscient settings where agents access complete information to maintain coherence. While this enhances performance, it fundamentally contradicts real-world social dynamics characterized by information asymmetry and private knowledge [87].
Relying on human evaluation to bridge this gap is perilous. Recent findings suggest that human raters are influenced by non-epistemic factors—such as the assertiveness and linguistic fluency of LLM outputs—leading to a potential overestimation of their factual accuracy [66]. This creates a ’validity illusion’: the simulation looks realistic because the text is fluent, but the stochastic and ”black-box” nature of LLMs makes calibration—the standard process of tuning ABM parameters to match observed outputs—notoriously difficult. Consequently, until quantitative calibration becomes a primary focus, large-scale LLM simulations remain in the realm of ’plausible storytelling’ rather than predictive science [44].
In light of these challenges, [81] proposes a pragmatic framework for determining the conditions under which LLM-based social simulations can meaningfully advance scientific understanding. They argue that these simulations are most defensible when used to explore collective patterns rather than individual decision trajectories, provided that agent outputs align with population-level averages and that validation methods are carefully matched to the research objectives and simulation scale.
4.2 The ”Black-Box” Problem and Interpretability
The ”black-box” nature of Large Language Models, stemming from architectures with billions of parameters and complex non-linear transformations, creates significant challenges for understanding their decision-making processes [47]. This opacity does not merely hinder interpretability; it actively contributes to ”automation bias.” This phenomenon manifests when modelers, seduced by the sophistication and apparent fluency of LLM outputs, over-attribute accuracy and authority to the model, even when its underlying reasoning is obscure or flawed [17].
This combination of technical opacity and misplaced human trust becomes dangerous in high-stakes social simulations. Because the model’s internal processing does not consistently mitigate biases [38], outputs can be uninterpretable, inaccurate, or discriminatory. When such simulations are used for critical applications—like policy testing, resource allocation, or urban planning—where direct human validation is often impossible [81], there is a substantial risk of propagating flawed conclusions. As noted by [47], regulators require assurances that decision-support tools are robust and accountable—requirements that the ”black-box” nature of LLMs directly contravenes. Without robust interpretability tools, the unexamined nature of these models could lead to misinformed policy decisions or exacerbate societal inequalities, even when the researchers’ intentions are positive.
Beyond these ethical concerns, using an LLM to generate agent behavior fundamentally alters the scientific modeling process itself. Unlike traditional Agent-Based Modeling (ABM), where the designer explicitly constructs behavioral rules based on theory or empirical evidence, LLM-driven agents bypass this stage. The behavioral model becomes implicit, embedded within billions of opaque parameters learned from massive datasets. As a result, the modeller relinquishes the ability to clearly define or justify the assumptions underlying agent behavior. This lack of epistemic control is especially problematic in scientific contexts, where the rationale for modeling choices must be transparent and subject to scrutiny. While LLMs may, in theory, reflect statistical patterns closer to real-world data, this advantage is undermined if the underlying mechanisms remain inaccessible and unvalidated.
4.3 Systemic Bias and the ”Average Persona” Risk
While Section 2.3 highlighted the intrinsic biases of LLMs, their impact scales non-linearly when applied to multi-agent populations. The most critical risk for social simulation is systemic homogenization. As noted by [81], the statistical nature of LLMs drives them to converge toward a generic, safe, and ”WEIRD” behavioral mean. In a simulation context, this creates a ”Replicant Effect”: even when agents are assigned distinct demographic prompts, their actual decision-making logic collapses into a uniform mode, effectively erasing minority behaviors and extreme viewpoints. Moreover, this homogenization is paradoxically accompanied by ”Generative Exaggeration”. As demonstrated by [52], when LLMs are prompted to adopt specific personas to escape the ”average,” they often drift into caricature, amplifying stereotypical traits to the point of toxicity rather than maintaining behavioral nuance. This suggests that current models struggle to find a middle ground between the ”flatness” of the average persona and the ”parody” of specific roles.
This ”Average Persona” problem is exacerbated by the ”Data Scarcity Wall.” As discussed regarding the limitations of fine-tuning, simulating specific, under-represented populations—such as smallholder farmers in the Global South—is often impossible because the massive, digitized behavioral logs required for training simply do not exist. Consequently, relying on generic LLMs to simulate these groups risks replacing missing data with hallucinated stereotypes.
To mitigate this, future research must move beyond mere prompt engineering. Developing more diverse and personalized training datasets is crucial, specifically incorporating first-person narratives, oral histories, and qualitative psychological insights to capture the ”long tail” of human behavior. Furthermore, incorporating external motivation structures (e.g., BDI architectures) can help force agents to deviate from the statistical mean by driving decisions through explicit goals and fears rather than mere text prediction. Without these structural interventions, simulations will remain valuable for exploring collective patterns of the ”average” human [81], but will fail to capture the diversity essential for understanding complex societal dynamics.
4.4 Computational Constraints and Scalability
While LLMs offer cost-efficiency compared to human experiments, they introduce a drastic computational bottleneck compared to traditional Agent-Based Models (ABMs). The resource intensity of LLMs—both in fine-tuning and real-time inference—limits practical scalability [16]. Crucially, this high cost is not merely a logistical hurdle but a methodological barrier: it hinders the application of classical sensitivity analyses, which require running simulations thousands of times to assess parameter influence and model robustness [56, 81]. Unlike lightweight rule-based agents that permit exhaustive parameter sweeps on standard hardware, the computational latency of LLMs currently forces researchers to trade statistical rigor for behavioral complexity, often preventing the extensive replications necessary to validate stochastic models.
4.5 Ethics in Application: Serious Games vs. Predictive Science
Given the methodological barriers outlined above, it is crucial to delineate the appropriate domains of application for LLM-based agents.
4.5.1 Interactive Environments and Serious Games
In contexts such as educational games and training simulations [71], LLMs offer direct and significant benefits. Here, the primary goal is often to create a ”believable” and ”engaging” experience rather than a perfectly verifiable replication of social reality [16]. LLMs excel at simulating dynamic, personalized interactions, enhancing user engagement through adaptive learning pathways and gamified narratives [30].
However, as highlighted by recent critiques, this application requires a shift from cautious optimism to critical vigilance. The primary risk lies in ’negative training transfer’: unlike traditional rule-based simulators where every outcome is verified, an LLM-based agent can hallucinate factual errors with high confidence. Recent studies highlight that this ’authoritative misinformation’ can lead to superficial learning and cognitive over-reliance [88]. Furthermore, instructional quality is threatened by inherent biases; as demonstrated by [76], models often perpetuate harmful stereotypes in personalized content. Consequently, the deployment of LLMs in training necessitates rigorous ’red-teaming’ and grounded architectures (RAG) to ensure the simulation remains a tool for learning, not confusion.
4.5.2 From Prediction to Exploratory Modeling
For ”pure social simulation” aimed at scientific inquiry, the implications of the ”validity gap” and ”epistemic opacity” (discussed in Sections 4.1 and 4.2) are decisive. Because of the risks of hallucination and the ”average persona” bias, we argue that LLMs are currently ill-suited for confirmatory empirical research or precise forecasting (e.g., predicting specific election outcomes or policy impacts).
Instead, their most scientifically robust application lies in Exploratory Modeling [14]. In this paradigm, LLMs serve not as oracles of truth, but as engines for hypothesis generation. They allow researchers to rapidly explore a vast space of possible social interactions and identify emergent patterns that—while not necessarily predictive of the real world—reveal the logical consequences of specific behavioral assumptions. As [81] suggests, these simulations should be viewed as ”reasoning pumps” to test collective dynamics under diverse narrative constraints, rather than as digital twins of reality.
5 Towards Hybrid Constitutional Architectures
The current landscape suggests a fundamental shift in the approach to social simulation. While LLMs offer compelling opportunities—enabling the rapid simulation of thousands of interactions and uncovering emergent behaviors [4, 56]—the limitations of ”off-the-shelf” models underscore the enduring value of traditional Agent-Based Models (ABMs).
5.1 Integrating LLMs into Classical Frameworks
The future of social simulation does not lie in ”off-the-world” AI agents, but in the seamless integration of LLMs within structured platforms such as GAMA [70] and NetLogo [34]. This approach, which we define as LLM-augmented ABM, addresses the fundamental problem of groundedness: while LLMs manipulate abstract concepts, the ABM provides the ”world model”—the physical, spatial, and resource-based constraints—that dictates the actual viability of AI-proposed actions.
This synergy allows researchers to move beyond purely narrative modeling. Recent work, such as Gao et al. (2025) on the S3 system, demonstrates how individual behavior alignment can be synchronized with complex social network structures, ensuring that LLM expressiveness does not sacrifice macroscopic coherence [26]. Simultaneously, frameworks like AgentTorch [15] open the door to massive scalability through GPU acceleration, enabling simulations of entire populations where the ABM manages the physics of interaction while the LLM simulates context-sensitive psychology.
For instance, in an epidemiological model like COMOKIT [27, 69], an LLM could interpret evolving social norms or disinformation, while the ABM maintains the structural rigor of disease spread. This hybrid necessity is even more acute in high-stakes environments. [82] recently demonstrated this in disaster evacuation simulations: while the ABM manages the physical constraints of crowd density and velocity to prevent unrealistic movements, the LLM-enhanced agents simulate the cognitive panic and decision-making under uncertainty. This ”Physics-Cognition” split ensures that the simulation remains physically valid while capturing the psychological volatility of human crowds.
5.2 Cognitive Alignment: Fine-Tuning and its Limits
To correct the instability of generic models, a promising path lies in behavioral alignment. Binz et al. (2025) have demonstrated that it is possible to transform an LLM into a cognitively aligned agent (the Centaur model) by training it on massive datasets derived from human psychological experiments. This process allows the AI to replicate human learning curves and decision-making biases correlated with human neural signals [10].
However, this data-driven path hits a critical ”Data Scarcity Wall” as soon as one moves beyond Western (WEIRD) contexts. As highlighted by [60], when representative training data is absent, LLMs tend to fall back on generalized stereotypes to fill the gaps. Consequently, in social simulation where identity is relational and cultural, modeling the decisions of smallholder farmers in Vietnam is impossible through simple fine-tuning: the necessary digitized behavioral logs do not exist, leaving the model to rely on hallucinated or westernized proxies.
Pursuing a strategy exclusively based on data risks reinforcing the ”Average Avatar Bias”: an AI that perfectly mimics a mean Western human but fails entirely to represent the specificities of marginalized or culturally distinct groups. To break this impasse, alignment must shift from being purely statistical (data-driven) to being structural (theory-driven), justifying the transition toward ”constitutional” architectures.
5.3 Proposed Paradigm: Hybrid Constitutional Architectures
We argue that cognitive alignment in social simulation cannot remain purely data-driven; it must be structurally anchored in psychological theory. We propose the adoption of Hybrid Constitutional Architectures, a paradigm shift where the LLM’s generative freedom is bounded by a theoretical ’scaffold.’
In this stratified cognitive process, we move beyond the monolithic agent:
  • Heuristic Stratum (SLMs): High-frequency, routine behaviors (e.g., daily navigation, basic consumption) are delegated to Small Language Models (SLMs). This ensures computational scalability [9] and allows for ’local alignment’—where smaller models are fine-tuned on specific ethnographic data that would be diluted in larger architectures.
  • Reasoning Stratum (LLM as Proposal Engine): Larger models are invoked only for complex, non-routine social reasoning. Crucially, the LLM does not execute the final decision; it acts as a ’Proposal Engine’, generating multiple plausible behavioral trajectories based on the current social context.
  • The Theoretical Constitution: Every proposal is passed through a validation filter grounded in established frameworks, such as the Theory of Planned Behavior (TPB) [2] or Bounded Rationality principles [65]. For instance, a proposal that contradicts an agent’s established ’Subjective Norms’ or ’Perceived Behavioral Control’ is rejected or penalized. The constitutional layer must itself be treated as a model component subject to audit, since poorly specified normative constraints may reduce harmful outputs while simultaneously suppressing legitimate social heterogeneity.
  • External Memory Modules (ABM State): To resolve the Temporal Instability inherent to finite context windows, the architecture offloads long-term state (biography, resources, social graph) to the classical ABM environment. Rather than relying on the LLM to ’remember’ the entire history, relevant data is retrieved dynamically via Retrieval-Augmented Generation (RAG). This separation ensures that agents maintain consistent identities over thousands of simulation steps [55, 54].
To translate this conceptual architecture into an explicit implementation protocol, we formalize the decision-making process as a stratified execution loop that strictly arbitrates between the ABM, SLMs, and LLMs. Let $E_{t}$ denote the global environment state at simulation step $t$, managed by the classical ABM (e.g., GAMA, NetLogo) acting as the absolute ground truth, and $S_{i,t}$ the internal state of agent $i$.
Arbitration is governed by a context-shift function, $\Delta(E_{t},S_{i,t})$, which continuously evaluates environmental and social stability. If $\Delta(E_{t},S_{i,t})\leq\tau$ (where $\tau$ is a predefined stability threshold), the environment is considered stable. The agent defaults to the computationally lightweight Heuristic Stratum, generating a routine action via a Small Language Model or classical heuristics: $a_{\text{heur}}=\text{SLM}(E_{t},S_{i,t})$.
However, if the ABM detects a novel event, a social anomaly, or a critical threshold breach (e.g., a sudden policy change or resource depletion), $\Delta(E_{t},S_{i,t})>\tau$. Decision-making is then escalated to the Reasoning Stratum. Here, the LLM acts as a proposal engine, generating an intended behavioral trajectory based on context and retrieved memory ($M_{i}$): $a_{\text{prop}}=\text{LLM}(E_{t},S_{i,t},M_{i})$.
Crucially, this generated intention is never executed directly. It must pass through the Theoretical Constitution, formalized as a validation filter $\Phi(a,E_{t},S_{i,t})\in\{0,1\}$. Grounded in psychological frameworks like the Theory of Planned Behavior (TPB), $\Phi$ evaluates both the physical possibility of the action within the ABM and its normative plausibility for the specific agent. The final action execution follows a strict fallback mechanism:
$a_{\text{final}}=\begin{cases}a_{\text{prop}}&\text{if }\Delta>\tau\text{ and }\Phi(a_{\text{prop}},E_{t},S_{i,t})=1\\ a_{\text{heur}}&\text{otherwise (fallback or routine)}\end{cases}$ (1)
Finally, the ABM gathers all validated actions ($\mathbf{A}$), resolves physical conflicts, and updates the environment state for the next step: $E_{t+1}=f_{\text{ABM}}(E_{t},\mathbf{A})$. This formal protocol (summarized in Algorithm 1) ensures that the generative power of LLMs is rigorously bounded by causal and computational constraints.
Algorithm 1 Hybrid Constitutional Execution Loop (Single Agent Step)
1:Current environment state $E_{t}$, Agent internal state $S_{i,t}$, Stability threshold $\tau$
2:Updated environment state $E_{t+1}$
[…]
5:if $\delta>\tau$ then $\triangleright$ Escalate to LLM Reasoning
6:  $M_{i}\leftarrow\text{RAG}(S_{i,t})$ $\triangleright$ Retrieve contextual memory
7:  $a_{\text{prop}}\leftarrow\text{LLM}(E_{t},S_{i,t},M_{i})$ $\triangleright$ Generate behavioral proposal
8:  if $\Phi(a_{\text{prop}},E_{t},S_{i,t})==1$ then $\triangleright$ Constitutional Validation
9:   $a_{\text{final}}\leftarrow a_{\text{prop}}$
10:  else
11:   $a_{\text{final}}\leftarrow a_{\text{heur}}$ $\triangleright$ Fallback due to constitutional rejection
12:  end if
13:else
14:  $a_{\text{final}}\leftarrow a_{\text{heur}}$ $\triangleright$ Default to routine behavior
15:end if
16:Execute $a_{\text{final}}$ via ABM engine
[…]
To measure the analytical improvement of this hybrid approach against both classical rule-based ABMs and pure LLM-driven simulations, we propose a tripartite evaluation pipeline. First, Counterfactual Robustness: unlike classical ABMs that fail outside pre-coded rules, and pure LLMs that may hallucinate physically impossible actions, the hybrid model must be tested against unseen socio-environmental shocks to ensure responses are both socially plausible and physically grounded. Second, Behavioral Diversity: since pure LLMs tend to collapse into an ’Average Persona,’ the hybrid architecture—particularly through locally fine-tuned SLMs—must be quantitatively evaluated on its ability to preserve statistical heterogeneity and minority behaviors. Finally, Computational Scalability: a key metric of success is the ability to run exhaustive sensitivity analyses (e.g., Monte Carlo simulations with 1000+ runs). While pure LLM models face prohibitive API latency and costs, the stratified delegation to SLMs must demonstrate a significant reduction in computational overhead, restoring the rigorous statistical validation expected in traditional social simulation.
By treating the LLM as a creative engine constrained by a formal psychological constitution, we solve the ’black-box’ problem of intent. Decision transparency is regained: the modeler can verify if a choice was made due to a social norm (rule-based) or a creative adaptation (LLM-based). This synergy moves the field from ’surface realism’ to causal plausibility, enabling simulations that remain robust even when empirical data is scarce. Rather than replacing classical agent-based modeling, Hybrid Constitutional Architectures reposition LLMs as bounded cognitive modules within a transparent causal framework, thereby reconciling expressive flexibility with epistemic accountability.
6 Conclusion
LLMs have demonstrated impressive capabilities in mimicking human-like conversation and passing complex psychological tests [35, 40]. Yet, as this paper has explored, these successes are often rooted in sophisticated pattern recognition rather than genuine understanding. This leads to the ”Fluency Fallacy”: a state where models maintain high linguistic coherence even when hallucinating social facts, creating a ”surface realism” that can mislead researchers into mistaking narrative plausibility for causal validity.
The transition from individual generative agents to robust social simulations requires moving beyond the Turing Test toward stricter criteria: Behavioral Diversity (to avoid mode collapse towards an average persona) and Counterfactual Robustness (the ability to react realistically to unseen shocks) [81, 44]. While LLMs show direct utility in interactive simulations and serious games for engagement [16], their application in pure social science demands a critical re-evaluation of epistemic control.
Ultimately, the future of social simulation does not lie in a choice between the data-driven flexibility of LLMs and the theory-driven rigor of ABM, but in their synergy. By coupling the narrative transparency of Generative AI with the structural rigor of Hybrid Constitutional Architectures, we can achieve simulations that are not only plausible but causally valid and intuitively explainable. This hybrid paradigm ensures that even in data-scarce environments, artificial societies can serve as meaningful laboratories for understanding the deep-seated nuances of human decision-making.
Conflict of Interest Statement
The authors declare that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.
Author Contributions
P.T. conceived the original structure and wrote the first draft of the manuscript. J.D.Z., A.G., B.G., N.Q.H., and A.D. substantially contributed to the conceptualization, provided critical feedback, and improved the manuscript. The development of this article was a truly collective effort. All authors discussed the concepts, read, and approved the final manuscript.